# First principle prediction of structural distortions in the cuprates and their impact on the electronic structure

### v2

@@ III.1 Crystal structure
For the oxygen-doped Bi2Sr2CaCu2222O8+x system with $x=0.25$, we account for the superlattice modulation Gao et al. (1988); Petricek et al. (1990); Levin et al. (1994); Slezak et al. (2008); He et al. (2008) and explore various possible positions for the oxygen dopants. These structural properties are essential for describing the BiO layers properly, and they also have a significant impact on the superconducting gap Andersen et al. (2007); Massee et al. (2020).
Prior work using smaller unit cells has shown that the lowest-energy oxygen dopant positions are between the Bi and Sr layers He et al. (2006); Nokelainen et al. (2020). Hence we follow the prior work to look for optimized oxygen dopant positions around the Bi and Sr layers. We have tested multiple oxygen dopant positions, and Fig. 4 (a) shows our optimized lowest-energy crystal structure for Bi2Sr2CaCu2O8.25, where four oxygen dopants are added into a stoichiometric undoped Bi-2212 unit cell of 240 atoms which is a $4\times 1\times 1$ enlargement of the 60-atom unit cell of Sec. II. Several metastable crystals are listed in the supplementary material Jin and Ismail-Beigi , showing an energy cost of at least 900 meV/dopant to move oxygen dopants between two BiO layers, and at least 168 meV/dopant to move dopants between BiO and SrO layers. Regardless of the dopant positions, the relaxed superlattice modulation remains in the same pattern, consistent with the modulations found in prior experimental and theoretical works Gao et al. (1988); Petricek et al. (1990); Levin et al. (1994); Slezak et al. (2008); He et al. (2008).
We conclude that the most stable positions for the oxygen dopants are located at the necking region between BiO and SrO layers as highlighted in Fig. 4. This position agrees with the in-plane dopant coordinates determined by STM experiments McElroy et al. (2005), but STM is performed on the surface and cannot describe the dopant positions along the out-of-plane direction. Hence, we compare directly to recent high-resolution STEM measurements Song et al. (2019) which can provide orthogonal information on atomic positions: we see excellent agreement with the STEM-observed structure as shown in Fig. 4 (b). Hence, we are confident in our predicted ground-state structure for the doped crystal and use it below to predict a variety of low-energy magnetic and charge orderings.
@@ III.2 Magnetic and charge ordering
With the lattice structure of our doped BSCCO confirmed, we turn our attention to the electronic spin and charge distribution within this system. We will show that many spatially spin- and/or charge-ordered electronic states are almost degenerate in energy with the G-type (checkerboard) AFM-ordered state. This indicates the presence of strong spin and charge fluctuations and competing orders in this overdoped normal state, and such fluctuations are thought to be a possible physical origin of the pseudogap phase in the cuprates Zhang et al. (2020).
We begin with short-period magnetic orderings of the Cu magnetic moments, such as the non-magnetic, ferromagnetic, G-AFM, and A-AFM states. Possible longer-period magnetic orders with spatial inhomogeneity, such as stripe order states, are considered further below. Not surprisingly, the most energetically favorable magnetic order among the short-period orderings is the G-AFM order with antiparallel nearest-neighbor spins on Cu atoms as illustrated in Fig. 5(a). Other short-period meta-stable magnetic orders exhibit aligned nearest-neighbor spins, either in an intralayer ($\sim 40$ meV/Cu higher in energy) or interlayer ($\sim 2$meV/Cu higher in energy) fashion, as discussed in the supplementary material Jin and Ismail-Beigi . Since the energy cost associated with changing the inter-bilayer spin alignment from antiparallel to parallel is negligible Jin and Ismail-Beigi , we will be concentrating below on the magnetic structure within a single bilayer.
@@
We have discovered numerous longer-ranged stripe-ordered states that exhibit nearly degenerate energies with the G-AFM state. While DFT-based stripe-ordered states were previously reported in LSCO Anisimov et al. (2004); Pesant and Côté (2011) and YBCO Zhang et al. (2020), they have not been observed in BSCCO up to now. For Bi-2212, we present a typical low-energy bond-centered stripe-ordered state in Fig. 5(b), where the nearest neighbor spins crossing the dashed domain walls align in parallel, in contrast to the antiparallel alignment in the G-AFM state. Remarkably, the total energy of this stripe-order state is only 1.9 meV/Cu higher than the G-AFM state. In Appendix B, we tabulate eight distinct stripe order patterns with energies with 3 meV/Cu of the G-AFM state. These stripe order states differ by their types and alignments of the domain walls, with only little energy differences. The existence of all these energetically competing orders suggests the presence of strong spin fluctuations in the normal state, which can play an important role in superconducting pairing Moriya (2006); Dahm et al. (2009); Wakimoto et al. (2004); Scalapino (2012). Figs. 5(c,d) show the local moment and $d_{x^{2}-y^{2}}$ occupancy of this stripe-order state. The sites closer to the domain walls have lower occupancies and smaller local moments than the other sites. These striped spin and charge orders present a modulation of ±$\pm 0.2\mu_{B}$ for local moments and ±$\pm 0.15$ for occupancies, significantly larger than the modulation caused by structural supermodulation captured in the G-AFM state. Consequently, the formation of the stripe order has an electronic origin, but the precise location of the domain boundaries can be influenced by the superlattice modulation effect.
@@
In addition to the Wannier basis described above, we have also analyzed the electronic structure of the stripe orders with the standard atomic projections output by the Vienna ab initio simulation package (VASP) software Kresse and Furthmüller (1996); Kresse and Furthmüller (1996) in the supplementary material Jin and Ismail-Beigi . The results are difficult to interpret due to the non-orthogonality of the standard VASP projections as explained in the supplement: the Cu magnetic moments show spatial modulations similar to those in Fig. 5 but the total occupancies hardly vary from site to site; additionally, there is a large oxygen contribution to the bands crossing the Fermi level. In short, in contrast to our $d_{x^{2}-y^{2}}$ Wannier basis and the intuitive picture it provides, the standard VASP projections are unable to explain the stripe order easily and we do not discuss them further. (The magnetic moments and electron counts for a three-band model are found in Appendix A which show qualitative agreement with our one-band model.)
Figure 6: Electronic structures of the 25% hole-doped Bi-2212 system with a non-magnetic state. (a) The unfolded band structure of the 244-atom unit cell. The opacity represents the spectral weight. (b) Schematic Fermi surfaces are black curves. The black, yellow, blue, and red rectangles represent the first BZ of the 1, 2, 4, and 8 Cu per layer unit cell, respectively. (c-e) Unfolded Fermi surfaces for the symmetrized crystal, the orthorhombic distorted crystal of Fig. 2(b), and the hole-doped crystal of Fig. 4, respectively. The Fermi level in (c) and (d) is shifted by 25% “virtual hole doping” to allow a fair comparison to the hole-doped Fermi surface in (e). Yellow dashed ($q_{1}$) and red solid ($q_{2}$) arrows show two different coupling wave vectors. (f) Measured ARPES Fermi surface at $T=104$ K, adapted from experiments with similar hole doping level $\sim 23\%$ He et al. (2021). Solid and dashed black curves highlight two distinct kinds of Fermi surface curves.
@@ III.3 ARPES spectra and shadow bands
The topology of the Fermi surface and the associated low-energy electronic spectrum, usually measured by ARPES Damascelli et al. (2003), provides important insights into the electronic properties of solids. In particular, many materials exhibit a so-called “shadow band” (SB) Fermi surface, resembling a weak-intensity copy of their main band (MB) Fermi surface with certain shifted vectors in momentum space. Depending on the system, these SB Fermi surfaces can, in principle, originate from any type of symmetry breaking such as electronic Brouet et al. (2008), magnetic Rotenberg et al. (2005), or structural Mans et al. (2006) origins. The physical origin of the SB Fermi surface is crucial for understanding the physical properties of the material, but it is difficult to distinguish between the different possible origins from ARPES measurements alone.
The Fermi surface of Bi-2212, as revealed by intensive ARPES studies, includes weaker intensity SB in addition to the main bands (see Fig. 6(f), adapted from an experiment He et al. (2021)). These SBs can be described by two types of symmetry-breaking vectors. One of them always aligns with the superstructural modulation direction, while the other is along ±$\pm(\pi,\pm\pi)$, coincident with the AFM ordering vectors. This has led to a continued debate regarding the magnetic Kampf and Schrieffer (1990); Langer et al. (1995); Chubukov (1995) versus the structural origin Koitzsch et al. (2004); Mans et al. (2006); Nakayama et al. (2006); Yu et al. (2020) of the ±$\pm(\pi,\pm\pi)$ folding vector. Direct theoretical interpretation of these SB Fermi surfaces using DFT has been lacking.
@@ III.3.1 Non-magnetic calculations
Due to the strong spin fluctuations suggested by the numerous competing stripe orders and the G-AFM states, the normal state of the hole-doped system should not be described by a single magnetically-ordered configuration (i.e., a single Slater determinant). In principle, an account of quantum spin fluctuations is needed for a comprehensive theoretical description including the magnetic susceptibility Stepanov et al. (2018) and spin correlations Parsons et al. (2016) in cuprates. In practice, however, there is an established recipe whereby spectra computed using band theory for a non-magnetic electronic state of the cuprates are nicely comparable to ARPES Fermi surface Damascelli et al. (2003); Sobota et al. (2021).
In this subsection, we use this pragmatic approach and consider the electronic band structure of the non-magnetic state to see what can be learned. The effect of spatial and temporal spin correlations and fluctuations on the computed ARPES spectra will be discussed in the following subsections. As an added benefit, removing the spin degree of freedom allows us to focus exclusively on the effects of symmetry breaking by structural perturbations. As demonstrated below, the non-magnetic state successfully accounts for many normal state spectroscopic properties, including the SB. To facilitate comparison to experimental ARPES spectra of the Fermi surface, we employ a standard “band-unfolding” method Ku et al. (2010); Brouet et al. (2012) to project the band structure of a large supercell onto the primitive unit cell Brillouin zone. This approach is known to reproduce spectral intensities observed in various ARPES experiments on many different materials qualitatively Lin et al. (2011); Medeiros et al. (2014); Tomić et al. (2014); Zhu et al. (2018). The band structures and unfoldings discussed below are computed using the one-band Wannier basis described above in Sec. III.2.
The panels of Fig. 6 display the unfolded band structures. Fig. 6(a) shows the case of the 244-atom 25% hole-doped system with a non-magnetic phase. Around the antinodal region near M/X’ in Fig. 6(a), we find two branches of flat bands below the Fermi level, which are split by the interlayer coupling in the bilayers, consistent with ARPES experiments He et al. (2021, 2018).
@@
Next, we turn to the effect of orthorhombic distortions on the Wannier tight-binding model given by the crystal in Fig. 2(b) without symmetrization. Similarly, 25% hole doping is introduced by lowering the Fermi level manually. This model exhibits an approximately circular-shape shadow band as shown in Fig. 6(d). This circular shape shadow band arises from the symmetry breaking of in-plane hoppings due to the orthorhombic distortions, as depicted in Fig. 2(b). Previous experiments have observed these shadow bands and suggested their likely structural origin Koitzsch et al. (2004); Mans et al. (2006); Nakayama et al. (2006) without providing a specific microscopic picture. We observe that the distortions in the Bi-O layers enlarge the unit cell and introduce an inter-band coupling at wave vectors $q_{1}=\pm(\pi,\pm\pi)$. This coupling leads to folding from the main bright curves to the shadow bands as indicated by the yellow dashed arrows in Fig. 6(d) and Fig. 6(e).
Finally, we study the Wannier model for the large supercell shown in Fig. 4 that includes both the Bi-O layer distortions and the superlattice modulation. Consequently, on top of the circle-like shadow band, there is another type of shadow band involving the wave vector $q_{2}=\pm(\pi/4,-\pi/4)$ shown in Fig. 6(e). This type of shadow band, often referred to as a “superstructure” in previous studies Liu et al. (2019), has long been attributed to a post-emission modulation effect from Bi-O layer buckling Yu et al. (2020). Prior experiments have also shown that this type of shadow band fades away when the crystal modulation is gradually reduced by Pb doping Liu et al. (2019). Our calculation presents a consistent and simple microscopic picture: the oxygen dopants and associated superlattice modulations further enlarge the unit cell and introduce an additional coupling at $q_{2}$. This coupling creates a further folding (illustrated by the red arrow) from the main bright curves to another set of shadow bands. Note that, in both experiments and our theory, this coupling only occurs in one crystalline direction due to the superlattice modulations being solely along the $a$-axis as shown in Fig. 4.
@@ III.3.2 Band theory treatment: static inhomogeneous local moments
DFT calculations (ours, as well as prior work, Zhang et al. (2020)) show that there are many competing low-energy states involving charge and spin ordering. Thus the simplest step going beyond the non-magnetic calculation is to stay with the DFT framework (i.e., band theory) and assume that the actual material has static local moments on each atomic site and is spatially inhomogeneous microscopically: different parts of the material have different local charge/spin orderings. This state is often known as the spin glass state. If experimental measurements average spatially over these inhomogeneities, one can attempt to describe the fact that. Like other cuprate superconductors, $x=0.25$ hole-doped BSCCO shows strong spin fluctuations Lyons et al. (1988); Brom and Zaanen (2003); Zhu et al. (2023); Lyons and Fleury (1988) and is paramagnetic (PM) Sachdev (2003); Yoshizaki et al. (1988); Kaiser et al. (2012) in experiments. Hence, the static spatial fluctuations of the charges/spins from the band theory should average to give a globally uniform and paramagnetic structure. The key point is that because band theory is based on a single Slater-determinant which precludes dynamic (temporal) electronic fluctuations, the fluctuations in the actual material must necessarily be described, within band theory, as static ones that are distributed inhomogeneously in space. In this section, we will see that this approach leads to poor or problematic predictions of ARPES spectra and that a method going beyond band theory will be needed for a consistent and accurate description.
@@
The simplest approach to dealing with inhomogeneity is to take a large library of DFT-computed low-energy structures and compute a thermally-averaged ARPES spectrum: for each spin/charge-ordered state, one computes the ARPES spectrum using band-unfolding and has it contribute to the average spectrum with a weight given by its thermal Boltzmann probability based on the energy of that state per unit cell. The physical picture is that each unit cell of the material is effectively independent of its neighboring cells so that the spatial inhomogeneity can be replaced by a thermal average over a single unit cell’s multiple possible states. In other words, one assumes that the domain size of each phase is on the order of a unit cell Markiewicz et al. (2019). We have performed this calculation using 8 different low-energy stripe order states as well as the G-AFM state for the 244-atom unit cell using a thermal energy of $k_{B}T=0.01$ eV. The resulting ARPES spectrum is shown in Fig. 7 and has little relation to the experimental observations.
Given this problematic spectral prediction, the above approach can be improved by relaxing the assumption of the fixed domain sizes: one should make a more faithful model that automatically chooses domain sizes and distributions based on total energy minimization in a large supercell, and then compute the associated ARPES spectrum of this large supercell that contains representative inhomogeneous charge/spin distributions Wang et al. (2020). It turns out (below) that one needs beyond 2,000 Cu sites in the supercell to converge the resulting ARPES spectra versus supercell size. Hence, this approach is not presently feasible using DFT due to its computational expense, and a surrogate model is needed. In the following, we will use a simple tight-binding approach based on Wannierization of the DFT band structure to arrive at a computationally tractable surrogate model.
@@
For BSCCO, we find that a high precision reproduction of the DFT band structure requires only three key types of hopping parameters which are $t=t_{\langle 100\rangle},t^{\prime}=t_{\langle 110\rangle}$ and $t^{\prime\prime}=t_{\langle 200\rangle}$ where $\langle 100\rangle$ is a nearest-neighbor hopping (along the Cu-O-Cu direction), etc. The Wannierization of the full 244-atom unit cell including the crystal modulation distortion yields us the hopping parameters $t_{ij}$ as well as the site-dependent $\epsilon_{i}$; the hopping parameters, when averaged over all appropriate pairs in the unit cell, have the values $t=-0.473$, $t^{\prime}=0.088$, and $t^{\prime\prime}=-0.091$ eV. (To simulate a high-symmetry structure, we symmetrize our model by setting all $\epsilon_{i}=0$ and all hoppings to these averaged values.) These hopping parameters are qualitatively consistent with the effective Hamiltonians in prior DFT studies on high-symmetry unrelaxed crystals Morée et al. (2022); Schmid et al. (2023); Ohgoe et al. (2020); Hirayama et al. (2019). In addition, due to the structural distortions and superlattice modulation, these hoppings are modulated spatially by about ±$\pm 0.015$eV, and the onsite energies are modulated by about ±$\pm 0.032$eV. The interaction parameter $U$ is chosen so that the HF total energy difference between the non-magnetic and GAFM phases at $x=0.25$ reproduces the total DFT energy difference between those phases. The resulting value is $U=3.1$ eV or $U/t\approx 7$, which is consistent with the effective Hamiltonian from prior works Morée et al. (2022); Schmid et al. (2023); Ohgoe et al. (2020); Nilsson et al. (2019); Hirayama et al. (2019), and is very reasonable compared to prior studies of cuprate systems using one-band Hubbard models Sheshadri et al. (2023); Dalla Piazza et al. (2012); Zheng et al. (2017).

### v3

@@ III.1 Crystal structure
For the oxygen-doped Bi2Sr2CaCu2222O8+x system with $x=0.25$, we explicitly include the superlattice modulation [34, 65, 66, 67, 68]. Experimentally, this modulation exists with a period between 4-5 unit cells regardless of doping level [92, 66, 93, 94], and theoretically confirmed that the modulation is an intrinsic behavior of BSCCO in that it occurs even without doping [68]. Of course, in a computational model using periodic boundary conditions, the superlattice modulation must conform to the size of the computational supercell, e.g., 5 unit cells in Ref. [68] and 4 unit cells in our work below; either choice is reasonable given the range of 4-5 unit cells found in experiments. We choose 4 unit cells for purely pragmatic reasons: the ARPES experiments [9] we will compare to below (concerning the shadow bands) are estimated to have a doping of $0.240.22<x<$ 0.24 (based on the superconducting temperature) or $x=0.26$ (based on the Fermi surface volume). Our choice of $x=0.25$ is suitable. Finally, unlike a calculation with periodic boundary conditions, the positions of the oxygen dopants may be disordered in the real material which will modify spectroscopic intensities. However, given the above-described robustness of the superlattice modulation with respect to doping, we believe that the superlattice modulation (and spectroscopic features deriving from it such as the shadow bands) should only suffer modest broadening and generally remain intact with respect to oxygen disorder.
For the chosen supercell which accommodates a period-4 superlattice, we explore various possible positions for the oxygen dopants. These structural properties are essential for describing the BiO layers properly, and they also have a significant impact on the superconducting gap [37, 95].
Prior work using smaller unit cells has shown that the lowest-energy oxygen dopant positions are between the Bi and Sr layers [76, 82]. Hence we follow the prior work to look for optimized oxygen dopant positions around the Bi and Sr layers. We have tested multiple oxygen dopant positions, and Fig. 4 (a) shows our optimized lowest-energy crystal structure for Bi2Sr2CaCu2O8.25, where four oxygen dopants are added into a stoichiometric undoped Bi-2212 unit cell of 240 atoms which is a $4\times 1\times 1$ enlargement of the 60-atom unit cell of Sec. II. Several metastable crystals are listed in the supplementary material [85], showing an energy cost of at least 900 meV/dopant to move oxygen dopants between two BiO layers, and at least 168 meV/dopant to move dopants between BiO and SrO layers. Regardless of the dopant positions, the relaxed superlattice modulation remains in the same pattern, consistent with the modulations found in prior experimental and theoretical works [34, 65, 66, 67, 68].
We conclude that the most stable positions for the oxygen dopants are located at the necking region between BiO and SrO layers as highlighted in Fig. 4. This position agrees with the in-plane dopant coordinates determined by STM experiments [36], but STM is performed on the surface and cannot describe the dopant positions along the out-of-plane direction. Hence, we compare directly to recent high-resolution STEM measurements [92] which can provide orthogonal information on atomic positions: we see excellent agreement with the STEM-observed structure as shown in Fig. 4 (b). Hence, we are confident in our predicted ground-state structure for the doped crystal and use it below to predict a variety of low-energy magnetic and charge orderings.
@@ III.2 Magnetic and charge ordering
With the lattice structure of our doped BSCCO confirmed, we turn our attention to the electronic spin and charge distribution within this system. We will show that many spatially spin- and/or charge-ordered electronic states are almost degenerate in energy with the G-type (checkerboard) AFM-ordered state. This indicates the presence of strong spin and charge fluctuations and competing orders in this overdoped normal state, and such fluctuations are thought to be a possible physical origin of the pseudogap phase in the cuprates [96].
We begin with short-period magnetic orderings of the Cu magnetic moments, such as the non-magnetic, ferromagnetic, G-AFM, and A-AFM states. Possible longer-period magnetic orders with spatial inhomogeneity, such as stripe order states, are considered further below. Not surprisingly, the most energetically favorable magnetic order among the short-period orderings is the G-AFM order with antiparallel nearest-neighbor spins on Cu atoms as illustrated in Fig. 5(a). Other short-period meta-stable magnetic orders exhibit aligned nearest-neighbor spins, either in an intralayer ($\sim 40$ meV/Cu higher in energy) or interlayer ($\sim 2$meV/Cu higher in energy) fashion, as discussed in the supplementary material [85]. Since the energy cost associated with changing the inter-bilayer spin alignment from antiparallel to parallel is negligible [85], we will be concentrating below on the magnetic structure within a single bilayer.
@@
We have discovered numerous longer-ranged stripe-ordered states that exhibit nearly degenerate energies with the G-AFM state. While DFT-based stripe-ordered states were previously reported in LSCO [102, 103] and YBCO [96], they have not been observed in BSCCO up to now. For Bi-2212, we present a typical low-energy bond-centered stripe-ordered state in Fig. 5(b), where the nearest neighbor spins crossing the dashed domain walls align in parallel, in contrast to the antiparallel alignment in the G-AFM state. Remarkably, the total energy of this stripe-order state is only 1.9 meV/Cu higher than the G-AFM state. In Appendix B, we tabulate eight distinct stripe order patterns with energies with 3 meV/Cu of the G-AFM state. These stripe order states differ by their types and alignments of the domain walls, with only little energy differences. The existence of all these energetically competing orders suggests the presence of strong spin fluctuations in the normal state, which can play an important role in superconducting pairing [104, 105, 106, 107]. Figs. 5(c,d) show the local moment and $d_{x^{2}-y^{2}}$ occupancy of this stripe-order state. The sites closer to the domain walls have lower occupancies and smaller local moments than the other sites. These striped spin and charge orders present a modulation of ±$\pm 0.2\mu_{B}$ for local moments and ±$\pm 0.15$ for occupancies, significantly larger than the modulation caused by structural supermodulation captured in the G-AFM state. Consequently, the formation of the stripe order has an electronic origin, but the precise location of the domain boundaries can be influenced by the superlattice modulation effect.
@@
In addition to the Wannier basis described above, we have also analyzed the electronic structure of the stripe orders with the standard atomic projections output by the Vienna ab initio simulation package (VASP) software [108, 109] in the supplementary material [85]. The results are difficult to interpret due to the non-orthogonality of the standard VASP projections as explained in the supplement: the Cu magnetic moments show spatial modulations similar to those in Fig. 5 but the total occupancies hardly vary from site to site; additionally, there is a large oxygen contribution to the bands crossing the Fermi level. In short, in contrast to our $d_{x^{2}-y^{2}}$ Wannier basis and the intuitive picture it provides, the standard VASP projections are unable to explain the stripe order easily and we do not discuss them further. (The magnetic moments and electron counts for a three-band model are found in Appendix A which show qualitative agreement with our one-band model.)
Figure 6: Electronic structures of the 25% hole-doped Bi-2212 system with a non-magnetic state. (a) The unfolded band structure of the 244-atom unit cell. The opacity represents the spectral weight. (b) Schematic Fermi surfaces are black curves. The black, yellow, blue, and red rectangles represent the first BZ of the 1, 2, 4, and 8 Cu per layer unit cell, respectively. (c-e) Unfolded Fermi surfaces for the symmetrized crystal, the orthorhombic distorted crystal of Fig. 2(b), and the hole-doped crystal of Fig. 4, respectively. The Fermi level in (c) and (d) is shifted by 25% “virtual hole doping” to allow a fair comparison to the hole-doped Fermi surface in (e). Yellow dashed ($q_{1}$) and red solid ($q_{2}$) arrows show two different coupling wave vectors. (f) Measured ARPES Fermi surface at $T=104$ K, adapted from experiments with similar hole doping level $\sim 23\%$ [9]. Solid and dashed black curves highlight two distinct kinds of Fermi surface curves.
@@ III.3 ARPES spectra and shadow bands
The topology of the Fermi surface and the associated low-energy electronic spectrum, usually measured by ARPES [110], provides important insights into the electronic properties of solids. In particular, many materials exhibit a so-called “shadow band” (SB) Fermi surface, resembling a weak-intensity copy of their main band (MB) Fermi surface with certain shifted vectors in momentum space. Depending on the system, these SB Fermi surfaces can, in principle, originate from any type of symmetry breaking such as electronic [111], magnetic [112], or structural [35] origins. The physical origin of the SB Fermi surface is crucial for understanding the physical properties of the material, but it is difficult to distinguish between the different possible origins from ARPES measurements alone.
The Fermi surface of Bi-2212, as revealed by intensive ARPES studies, includes weaker intensity SB in addition to the main bands (see Fig. 6(f), adapted from an experiment [9]). These SBs can be described by two types of symmetry-breaking vectors. One of them always aligns with the superstructural modulation direction, while the other is along ±$\pm(\pi,\pm\pi)$, coincident with the AFM ordering vectors. This has led to a continued debate regarding the magnetic [113, 114, 115] versus the structural origin [116, 35, 117, 118] of the ±$\pm(\pi,\pm\pi)$ folding vector. Direct theoretical interpretation of these SB Fermi surfaces using DFT has been lacking.
@@ III.3.1 Non-magnetic calculations
Due to the strong spin fluctuations suggested by the numerous competing stripe orders and the G-AFM states, the normal state of the hole-doped system should not be described by a single magnetically-ordered configuration (i.e., a single Slater determinant). In principle, an account of quantum spin fluctuations is needed for a comprehensive theoretical description including the magnetic susceptibility [119] and spin correlations [120] in cuprates. In practice, however, there is an established recipe whereby spectra computed using band theory for a non-magnetic electronic state of the cuprates are nicely comparable to ARPES Fermi surface [110, 121].
In this subsection, we use this pragmatic approach and consider the electronic band structure of the non-magnetic state to see what can be learned. The effect of spatial and temporal spin correlations and fluctuations on the computed ARPES spectra will be discussed in the following subsections. As an added benefit, removing the spin degree of freedom allows us to focus exclusively on the effects of symmetry breaking by structural perturbations. As demonstrated below, the non-magnetic state successfully accounts for many normal state spectroscopic properties, including the SB. To facilitate comparison to experimental ARPES spectra of the Fermi surface, we employ a standard “band-unfolding” method [122, 123] to project the band structure of a large supercell onto the primitive unit cell Brillouin zone. This approach is known to reproduce spectral intensities observed in various ARPES experiments on many different materials qualitatively [124, 125, 126, 127]. The band structures and unfoldings discussed below are computed using the one-band Wannier basis described above in Sec. III.2.
The panels of Fig. 6 display the unfolded band structures. Fig. 6(a) shows the case of the 244-atom 25% hole-doped system with a non-magnetic phase. Around the antinodal region near M/X’ in Fig. 6(a), we find two branches of flat bands below the Fermi level, which are split by the interlayer coupling in the bilayers, consistent with ARPES experiments [9, 128].
@@
Next, we turn to the effect of orthorhombic distortions on the Wannier tight-binding model given by the crystal in Fig. 2(b) without symmetrization. Similarly, 25% hole doping is introduced by lowering the Fermi level manually. This model exhibits an approximately circular-shape shadow band as shown in Fig. 6(d). This circular shape shadow band arises from the symmetry breaking of in-plane hoppings due to the orthorhombic distortions, as depicted in Fig. 2(b). Previous experiments have observed these shadow bands and suggested their likely structural origin [116, 35, 117] without providing a specific microscopic picture. We observe that the distortions in the Bi-O layers enlarge the unit cell and introduce an inter-band coupling at wave vectors $q_{1}=\pm(\pi,\pm\pi)$. This coupling leads to folding from the main bright curves to the shadow bands as indicated by the yellow dashed arrows in Fig. 6(d) and Fig. 6(e).
Finally, we study the Wannier model for the large supercell shown in Fig. 4 that includes both the Bi-O layer distortions and the superlattice modulation. Consequently, on top of the circle-like shadow band, there is another type of shadow band involving the wave vector $q_{2}=\pm(\pi/4,-\pi/4)$ shown in Fig. 6(e). This type of shadow band, often referred to as a “superstructure” in previous studies [129], has long been attributed to a post-emission modulation effect from Bi-O layer buckling [118]. Prior experiments have also shown that this type of shadow band fades away when the crystal modulation is gradually reduced by Pb doping [129]. Our calculation presents a consistent and simple microscopic picture: the oxygen dopants and associated superlattice modulations further enlarge the unit cell and introduce an additional coupling at $q_{2}$. This coupling creates a further folding (illustrated by the red arrow) from the main bright curves to another set of shadow bands. Note that, in both experiments and our theory, this coupling only occurs in one crystalline direction due to the superlattice modulations being solely along the $a$-axis as shown in Fig. 4.
@@ III.3.2 Band theory treatment: static inhomogeneous local moments
DFT calculations (ours, as well as prior work, [96]) show that there are many competing low-energy states involving charge and spin ordering. Thus the simplest step going beyond the non-magnetic calculation is to stay with the DFT framework (i.e., band theory) and assume that the actual material has static local moments on each atomic site and is spatially inhomogeneous microscopically: different parts of the material have different local charge/spin orderings. This state is often known as the spin glass state. If experimental measurements average spatially over these inhomogeneities, one can attempt to describe the fact that. Like other cuprate superconductors, $x=0.25$ hole-doped BSCCO shows strong spin fluctuations [130, 131, 132, 133] and is paramagnetic (PM) [134, 135, 136] in experiments. Hence, the static spatial fluctuations of the charges/spins from the band theory should average to give a globally uniform and paramagnetic structure. The key point is that because band theory is based on a single Slater-determinant which precludes dynamic (temporal) electronic fluctuations, the fluctuations in the actual material must necessarily be described, within band theory, as static ones that are distributed inhomogeneously in space. In this section, we will see that this approach leads to poor or problematic predictions of ARPES spectra and that a method going beyond band theory will be needed for a consistent and accurate description.
@@
The simplest approach to dealing with inhomogeneity is to take a large library of DFT-computed low-energy structures and compute a thermally-averaged ARPES spectrum: for each spin/charge-ordered state, one computes the ARPES spectrum using band-unfolding and has it contribute to the average spectrum with a weight given by its thermal Boltzmann probability based on the energy of that state per unit cell. The physical picture is that each unit cell of the material is effectively independent of its neighboring cells so that the spatial inhomogeneity can be replaced by a thermal average over a single unit cell’s multiple possible states. In other words, one assumes that the domain size of each phase is on the order of a unit cell [137]. We have performed this calculation using 8 different low-energy stripe order states as well as the G-AFM state for the 244-atom unit cell using a thermal energy of $k_{B}T=0.01$ eV. The resulting ARPES spectrum is shown in Fig. 7 and has little relation to the experimental observations.
Given this problematic spectral prediction, the above approach can be improved by relaxing the assumption of the fixed domain sizes: one should make a more faithful model that automatically chooses domain sizes and distributions based on total energy minimization in a large supercell, and then compute the associated ARPES spectrum of this large supercell that contains representative inhomogeneous charge/spin distributions [138]. It turns out (below) that one needs beyond 2,000 Cu sites in the supercell to converge the resulting ARPES spectra versus supercell size. Hence, this approach is not presently feasible using DFT due to its computational expense, and a surrogate model is needed. In the following, we will use a simple tight-binding approach based on Wannierization of the DFT band structure to arrive at a computationally tractable surrogate model.
@@
For BSCCO, we find that a high precision reproduction of the DFT band structure requires only three key types of hopping parameters which are $t=t_{\langle 100\rangle},t^{\prime}=t_{\langle 110\rangle}$ and $t^{\prime\prime}=t_{\langle 200\rangle}$ where $\langle 100\rangle$ is a nearest-neighbor hopping (along the Cu-O-Cu direction), etc. The Wannierization of the full 244-atom unit cell including the crystal modulation distortion yields us the hopping parameters $t_{ij}$ as well as the site-dependent $\epsilon_{i}$; the hopping parameters, when averaged over all appropriate pairs in the unit cell, have the values $t=-0.473$, $t^{\prime}=0.088$, and $t^{\prime\prime}=-0.091$ eV. (To simulate a high-symmetry structure, we symmetrize our model by setting all $\epsilon_{i}=0$ and all hoppings to these averaged values.) These hopping parameters are qualitatively consistent with the effective Hamiltonians in prior DFT studies on high-symmetry unrelaxed crystals [59, 60, 55, 54]. In addition, due to the structural distortions and superlattice modulation, these hoppings are modulated spatially by about ±$\pm 0.015$eV, and the onsite energies are modulated by about ±$\pm 0.032$eV. The interaction parameter $U$ is chosen so that the HF total energy difference between the non-magnetic and GAFM phases at $x=0.25$ reproduces the total DFT energy difference between those phases. The resulting value is $U=3.1$ eV or $U/t\approx 7$, which is consistent with the effective Hamiltonian from prior works [59, 60, 55, 56, 54], and is very reasonable compared to prior studies of cuprate systems using one-band Hubbard models [139, 140, 100].